# Accurate Low Rank Approximation at a Low Computational Cost

### v13

@@ 6 Numerical experiments
We devote the next three subsections to the tests of our LRA algorithms on synthetic and real-world input matrices devote Sec. 6.4 to testing Alg. 5.1.
We implemented our algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory. We called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
@@ 6.1 Input Matrices
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [20, Sec. 7.1] for more details).
We display the distribution of the top 50 singular values of these three matrices in Fig. 1.
@@
Figure 1: Spectra of singular values of some real world matrices
Furthermore, we computed rank-$r$ approximations for the three matrices, StreamVel, MinTemp, and MaxCut, used in numerical experiments in [55, Sec. 7.3.2].
@@
In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2 and set the target rank $r$ equal to 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
@@
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the next three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [55, Sec. 7.3.1], which involved effective rank parameter $R=20$.
(i) Low-rank + noise matrices
${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$
for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, also involves parameter $\xi$, equal to $10^{-4}$ for the Low Rank Low Noise matrix, to $10^{-2}$ for the Low Rank Med Noise matrix, and to $10^{-1}$ for the Low Rank High Noise matrix.
(ii) We generated Polynomial Decay matrices
${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p}),$
where $p=0.5$ for Poly Decay Slow, $p=1$ for Poly Decay Med, $p=2$ for Poly Decay Fast.
(iii) We generated Exponential Decay matrices
${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q}),$
with $q=0.01$ for Exp Decay Slow, $q=0.1$ for Exp Decay Med, $q=0.5$ for Exp Decay Fast.
6.2 Two options for Alg. $\mathbb{A}$
In our tests of Algs. 2.1 and 3.1 we used the following two options for Alg. $\mathbb{A}$, which computes a crude rank-$\rho$ approximation of a matrix $M$.
(i) Fast (although not superfast) randomized algorithm of [54] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times(\rho)$.
(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{R}^{(2\rho)\times 1024}$ and $H\in\mathbb{R}^{1024\times\rho}$ replaced Gaussian ones.
6.3 Test results for Algs. 2.1 and 3.1
In our tests of escalation Alg. 2.1 we output the relative error norms
$\text{err}=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$
@@
In our tests of iterative escalation Alg. 3.1 we output the relative error norms
[…]
We only display the error norms output in the first three iterations of Alg. 3.1 and notice that, except for the matrices representing high noise or slow decay, the second iteration greatly decreased the output error norms, while the third iteration tended to keep the error norms at more or less the same level.
[…]
For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple of $r$, $\rho$, and an input matrix type for Alg. 2.1, and per the pair of $r$ and an input matrix type for Alg. 3.1, except that we ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger size.
@@
In Tables 6.1, 6.2, 6.5, and 6.6 we display the results of our tests with both Gaussian random and abridged SRHT test matrices. This covers the cases where the output error norms were more or less in the same range for both test matrices (and usually quite similar in both cases, with only a slightly better results with Gaussian random test matrices).
Displaying the results of our tests with the synthetic input matrices from [55, Sec. 7.3.1] in Tables 6.3 and 6.4, however, we only show the test results with Gaussian random test matrices because all tests with abridged SRHT test matrices have failed, that is, have output with large errors. The tests with Gaussian random test matrices have output accurate LRAs for the same input matrices, except for those representing high noise or slow decay.
@@
In both cases of success and failure of both algorithms, our test results were in good accordance with our formal study. The relative error norms output by both Algs. 2.1 and 3.1 have consistently decreased and frequently to the optimal value 1 (up to within the measurement errors) as the spectra of the input matrices decayed. The error norms have also decreased as we increased the upper rank $\rho$ for Alg. 2.1 and, as we said already, in the transition from the first to the second iteration of Alg. 3.1.
The error norm bounds of Alg. 2.1 for $\rho=4r$ were consistently smaller than those of Alg. 3.1 in all three iterations, except for tests with the matrix Shaw for $r=20$. This exception implies that our estimate of Cor. 3.2 is overly pessimistic for worst case inputs, as we explained in Remark 3.2.
@@
Table 6.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, synthetic matrices with fast decaying spectra (Fast Decay), and synthetic matrices with slow decaying spectra (Slow Decay).
@@
Table 6.2: Test results for Alg. 3.1 for $\rho=2r$ on synthetic matrices with fast decaying spectra (Fast Decay), synthetic matrices with slow decaying spectra (Slow Decay), Shaw, Gravity, and discretized single layer potential operator (SLP).
@@
Table 6.3: Test results for Alg. 2.1 on the synthetic input matrices from [55, Sec. 7.3.1] using Gaussian multipliers.
@@
Table 6.4: Test results for Alg. 3.1 on the synthetic input matrices from [55, Sec. 7.3.1] using Gaussian multipliers.
@@
Table 6.5: Test results for Alg. 2.1 on the real world input matrices from [55, Sec. 7.3.2]
@@
Table 6.6: Test results for Alg. 3.1 for $\rho=2r$ on real world input matrices from [55, Sec. 7.3.2].
6.4 Test results for Alg. 5.1
For each $n\times n$ matrix $M$, we computed its rank-10 approximation $\tilde{M}$ by its truncated SVD and then obtained the error matrix $E=M-\tilde{M}$. We generated initial vectors ${\bf v}$ according to the following
INITIALIZATION RULE: Define a vector ${\bf v}$ of length $n$ with coordinates $1/k$ at $k$ positions sampled at random independently under the uniform probability distribution; fill the vector with 0s elsewhere; do this for $k=1,\log\log(n),\log(n)\text{ and }n$.
We performed 100 tests for each pair of input matrix type and integer $k$ and measured the accuracy of the estimates by the ratio
$\text{accuracy}=\frac{||E||_{1}}{\text{est}},$
where est denoted the estimate of $||E||_{1}$ returned by the algorithm.
The ratio took on its minimal value 1 where the estimate est was optimal.
In our test most of the norm estimates est returned by the algorithm were within a factor of 2 from the optimal norm $||E||_{1}$, even where the initial vector had just a single nonzero entry.
All our tests converged within 6 iterations, regardless of the density of the initial vector.
@@ 7 Conclusions
We proposed, tested, and analyzed some simple but novel LRA algorithms using escalation and proved that at a dominated or even sublinear additional computational cost they significantly decrease the spectral norms of the output error matrices of the known algorithms applied to FDS matrices.
The results of our numerical tests with a variety of synthetic and real world matrices are in good accordance with our formal estimates.
Furthermore, some superfast (sublinear cost) variants of the known algorithms are known to output LRA that tend to be less accurate. In combination with our algorithms, however, this deficiency has been fixed in most of our tests, and it is a challenge to test the output accuracy of these and other superfast (sublinear cost) LRA algorithms for a greater variety of inputs.
Towards enhancing our progress to a larger class of matrices $M$, we nontrivially extended to LRA the popular technique of iterative refinement. So far the resulting algorithm was superior to our escalation LRA for only a single input class of matrices among 17 classes tested, but further comparative study may still be in order.
Another natural but yet unexplored challenge is to compare and possibly combine the iterative refinement LRA algorithm of [25] with our present escalation and iterative refinement algorithms.
@@ Appendix A Small families of matrices that are hard for superfast LRA
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Furthermore, the same argument shows that superfast estimation of the norm of a matrix fails already on the same matrix family and that for both problems of LRA and norm estimation randomized superfast algorithms fail with error probability not close to 0.
@@ Appendix B Computation of $r$-top SVD of a matrix product
For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [31, 32, 37, 46, 38]) based on computation of the SVDs of $A$ and $B$.
@@
Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{R}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
@@
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [15, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
@@ C.1 CUR decomposition and CUR approximation
of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$ (cf. [8]), although we recall notation $G$ for nucleus and CGR for the factorization in [30], originated with the decomposition itself in the 1990s [17]. In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
@@
If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large, but empirically the Cross-Approximation (C-A) celebrated iterations, briefly recalled in Appendix E, tend to converge fast to quite an accurate rank-$r$ CUR approximation for a large class of matrices $M$.
@@ C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{R}^{m\times n}$, $\Sigma\in\mathbb{R}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{R}^{m\times r}$, and $V\in\mathbb{R}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
@@
Three matrices $C\in\mathbb{R}^{m\times r}$, $N\in\mathbb{R}^{r\times r}$, and $R\in\mathbb{R}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
@@ Appendix D Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in Alg. 1.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [26, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
@@
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [20, Sec. 11]).
In this Appendix we follow [40, 41, 43, 42] where abridged SRFT matrices have been defined and tested similarly to abridged SRHT matrices with similar test results.
Appendix E Cross-Approximation (C-A) iterations
In this appendix we briefly recall C-A iterations, which can be viewed as a specialization to LRA of Alternative Directions Implicit (ADI) method.
C-A iterations:
  • For an $m\times n$ matrix $M$ and a target rank $r$, fix four integers $k$, $\ell$, $p$ and $q$ such that $0<r\leq k\leq p\leq m$ and $r\leq\ell\leq q\leq n$. [C-A iterations are simplified if $p:=k$ and $q:=\ell$.]
  • Fix an $m\times q$ “vertical” submatrix of the matrix $M$, made up of its $q$ columns. [^14] Of course, one can alternatively begin C-A iterations with a “horizontal” submatrix.
  • By applying a fixed CUR LRA subalgorithm, e.g., one of the algorithms of [29, 14, 33, 8], [^15] Such a subalgorithm runs at sublinear cost on the inputs of smaller size involved in C-A iterations, even where the original algorithms of [29, 14, 33, 8] applied to an $m\times n$ matrix $M$ run at superlinear cost. compute a $k\times\ell$ CUR generator $G$ of this submatrix and reuse it for the matrix $M$.
  • Compute the rank-$r$ CUR of $M$ based on the generator $G$ and output this CUR it if it is close enough.
  • Otherwise swap $p$ and $q$ and reapply the algorithm to the matrix $M^{*}$.
    [This is equivalent to computing a $k\times\ell$ CUR generator of a fixed $p\times n$ “horizontal” submatrix $M_{1}$ of $M$ that covers the submatrix $G$.]
  • Recursively alternate such “vertical” and “horizontal” steps until an accurate CUR LRA is computed or until the number of recursive C-A steps exceeds a fixed tolerance bound.
Figure 2: The first three recursive C-A steps output three striped submatrices.
Initialization recipes. For a large class of input matrices it is efficient to apply adaptive C-A iterations (cf. [1, 2]). They combine initialization with dynamic search for gaps in the singular values of $M$ by properly adapting Gaussian elimination with pivoting. The alternative initialization in [16] uses $O(nr^{2})$ flops to initialize the C-A iterations for an $n\times n$ input and $q=s=k=\ell=r$, and then the algorithm uses $O(rn)$ flops per C-A step. The output of C-A iterations, however, can be off by a significant factor.
Appendix F Iterative refinement of LRA with no compression
The definition of $r$-truncation immediately implies
Lemma F.1.
$M_{s}+(M-M_{s})_{r}=M_{r+s}$ for any pair of positive integers $r$ and $s$.
Next, based on this lemma, we lift $r$-top SVD into $kr$-top SVD by applying iterative refinement of LRA with no compression.
Algorithm F.1.
[Lifting $r$-top SVD into $kr$-top SVD.]
Input:
An $m\times n$ matrix $M$, $n\leq m$, two positive integers $k$ and $r$, an Alg. $\mathbb{S}$ that adds or subtracts a pair of $m\times n$ matrices, and an Alg. $\mathbb{T}$ that computes the $r$-truncation of an $m\times n$ matrix.
Output:
The $kr$-truncation of $M$.
Initialization:
Write $X_{0}:=0$ and $E_{0}:=M$.
Computations:
Recursively combine Algs. $\mathbb{S}$ and $\mathbb{T}$ to compute the matrices
$X_{i+1}:=X_{i}+(E_{i})_{r},~{}E_{i+1}:=M-X_{i+1},~{}~{}i=0,1,\dots,k-1.$ (F.1)
Output the matrix $X_{k}$.
The following proposition implies correctness of the algorithm.
Proposition F.1.
It holds that $X_{i}=M_{ir}$ for the matrices M and $X_{i}$ of Alg. F.1 for $i=0,1,\dots,k$.
Proof.
The claim holds for $i=0$ because $M_{0}=0$. Let it hold for $i\leq j$, that is, let $X_{j}=M_{jr}$, and then deduce from Eqn. (F.1) for $i=j$ that $X_{j+1}=X_{j}+(M-X_{j})_{r}$. By assumption, $X_{j}=M_{jr}$. Combine the latter equations and obtain $X_{j+1}=M_{jr}+(M-M_{jr})_{r}$. Apply Lemma F.1 for $s=jr$ and obtain that $X_{j+1}=M_{(j+1)r}$. Now the proposition follows by induction. ∎
Trying to extend our error analysis of Alg. 3.1 to Alg. F.1 we needed to estimate for the norm of the perturbation of the $r$-top SVD, but at this point the known studies (cf. [15, Thm. 8.6.5], [52, Thm. 3.3.4]) provided much weaker results than Thm. 1.2, and we ended with no support for numerical stability of Alg. F.1.

### v16

@@ 5 Numerical experiments
In this section we cover the tests of our LRA algorithms on synthetic and real-world matrices. We implemented the algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory and called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
@@ 5.1 Input Matrices
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [11, Sec. 7.1] for more details).
Fig. 1 shows the distribution of the top 50 singular values of these three matrices.
@@
Figure 1: Spectra of singular values of some real-world matrices
We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [34, Sec. 7.3.2].
@@
We let the target rank $r$ be 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2.
@@
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [34, Sec. 7.3.1], which involved effective rank parameter $R=20$.
(i) Low-rank + noise matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$ for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, depend on a parameter $\xi$ and become Low Rank Low Noise matrix for $\xi=10^{-4}$, the Low Rank Med Noise matrix for $\xi=10^{-2}$, and the Low Rank High Noise matrix for $\xi=10^{-1}$.
(ii) Polynomial Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p})$ are Poly Decay
Slow
for $p=0.5$, Poly Decay Med for $p=1$, and Poly Decay Fast for $p=2$.
(iii) Exponential Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q})$ are Exp Decay
Slow
for $q=0.01$, Exp Decay Med for $q=0.1$, and Exp Decay Fast for $q=0.5$.
5.2 Two options at stages 1 of Algs. 2.1 and 4.1
In our tests of Algs. 2.1 and 4.1 we used the following two options at their stage 1 for computations of a crude rank-$\rho$ approximation of a matrix $M$.
(i) Fast (although not superfast) randomized algorithm of [33] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{C}^{(2\rho)\times 1024}$ and $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
5.3 Test results for Algs. 2.1 and 4.1
For Alg. 2.1 we output the relative error norms $\text{err}:=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$ for input matrices $M$, a target rank $r$, the $r$-truncation $M_{r}$ and $M(\rho)_{r}$, and the upper rank $\rho$, set to $2r,3r,4r$, $5r$ (cf. Remark 2.3).
@@
For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg. 2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg. 4.1 where TYPE stands for an input matrix type. We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT test matrices, which were more or less in the same range for both matrix types. All our tests with the synthetic input matrices from [34, Sec. 7.3.1] wit abridged SRHT test matrices have failed, that is, have output large error norms. In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random test matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
@@
In both cases of success and failure of both algorithms, our test results were in good accordance with our formal study. The relative error norms output by both Algs. 2.1 and 4.1 applied to $(r,\rho)$-matrices have consistently decreased and frequently to the optimal value 1 (up to the measurement errors) as we increased the upper rank $\rho$ for Alg. 2.1 and in the transition from the first to the second iteration of Alg. 4.1. In all tests except for High Noise, Slow Decay, and MinTemp matrices already err2 decreased below 1.1. The third iteration decreased the error norm much less or even slightly increased it.
For $(r,\rho)$-matrices the error norm bounds of Alg. 2.1 for $\rho=4r$ were consistently smaller than those of Alg. 4.1 in all three iterations. This was not the case just for the matrices with flat spectra such as the matrices Shaw for $r=20$, High Noise, and Exp Decay Slow.
@@
Table 5.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, and synthetic matrices Fast Decay and Slow Decay, with fast and slowly decaying spectra.
@@
Table 5.2: Test results for Alg. 4.1 on Fast Decay, Slow Decay, Shaw, Gravity, and SLP matrices.
@@
Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian test matrices.
@@
Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian multipliers.
@@
Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [34, Sec. 7.3.2]
@@
Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [34, Sec. 7.3.2].
@@ 6 Conclusions
Rank-$r$ approximation of an $m\times n$ matrix $M$ is only interesting if there is a gap in the spectrum of its top (largest) $r+1$ singular values $\sigma_{j}(M)$. For the $(r,\rho)$-matrices $M$ having also a gap in the next segment of their spectra – from $(r+1)$st to $(\rho+1)$st largest singular values – randomized oversampling algorithms output quite accurate LRA.
We extended these observations towards devising superfast (sublinear cost) variation of LRA algorithms based on random sampling and oversampling with sparse test matrices and to improving accuracy of the resulting LRA in the case of $(r,\rho)$-matrices $M$ where $r<\rho\ll\sqrt{n}$. We provided some formal and experimental support for the efficiency of the resulting LRA algorithms.
There are various natural extensions of our study, which can be practically competitive. E.g., one can devise and test similar variations of the TYUC and GN algorithms with sparse test matrices such as Ultra-Sparse Rademacher test matrices [33, Sec. 3.9] or sparsified random Givens rotations by Rokhlin and Tygert (see [11, Remark 4.6]).666[26, 27, 28] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes.
Our study in Secs. 4 and 5 has showed no benefits for using iterative refinement of LRA versus oversampling so far, but one can keep comparing their efficiency.
@@ Appendix A Small families of matrices that are hard for superfast LRA
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Likewise, we can verify that any superfast randomized LRA algorithms fail fails on those matrix families with error probability not close to 0.
@@ Appendix B Computation of $r$-top SVD of a matrix product
For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [22, 30]) based on computation of the SVDs of $A$ and $B$.
@@
Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{C}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
@@
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [7, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
@@ C.1 CUR decomposition and CUR approximation
of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$, although we recall notation $G$ for nucleus and CGR for the factorization in [21], originated with the decomposition itself in the 1990s [8]. In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
@@
If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large.
@@ C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{C}^{m\times n}$, $\Sigma\in\mathbb{C}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{C}^{m\times r}$, and $V\in\mathbb{C}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
@@
Three matrices $C\in\mathbb{C}^{m\times r}$, $N\in\mathbb{C}^{r\times r}$, and $R\in\mathbb{C}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
@@ Appendix D Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in our tests for Algs. 2.1 and 4.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [17, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
@@
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [11, Sec. 11]).
Acknowledgement: Jianlin Xia helped place our study into proper historical context.